# Algorithms for Optimizing Acyclic Queries

### v2

@@ 4 Enumerating Join Trees
Our strategy for enumerating join trees is based on the following classic result relating join trees of a hypergraph to maximum spanning trees (MSTs) of its line graph:
Proposition 6 (Maier [27]).
A subgraph of $L(H)$ is a join tree of $H$ if and only if it is a maximum spanning tree of $L(H)$ according to the weight function $\omega$.
A naïve approach is to apply an off-the-shelf algorithm for enumerating the maximum spanning trees of $L(H)$. The best known algorithm for this purpose is due to Eppstein [12], and it works by deriving from the input graph $G$ a so-called equivalent graph $G^{\equiv}$. Every spanning tree of $G^{\equiv}$ corresponds to an MST of $G$ and vice versa. Eppstein proves a lower bound of $\Omega(m+n\log n)$ on constructing the equivalent graph from an arbitrary weighted graph with $m$ edges and $n$ vertices. Then to enumerate all $k$ MSTs of $G$, Eppstein applies existing algorithms to enumerate the spanning trees of $G^{\equiv}$. Since there are optimal spanning tree enumeration algorithms that run in $\mathcal{O}(m+n+k)$ [22, 35], the overall time complexity to enumerate MSTs is $\mathcal{O}(m+n\log n+k)$.
The main result of this section is an algorithm for enumerating join trees leveraging the structure of acyclic hypergraphs and their line graphs. In particular:
  • for any $\alpha$-acyclic hypergraph $H$, we can construct an equivalent graph in $\mathcal{O}(|L|)=\mathcal{O}(m+n)$, where $m$ and $n$ are the numbers of edges and vertices in $L(H)$, thus enumerating the join trees in $\mathcal{O}(|L|+k)$.
  • for any $\gamma$-acyclic hypergraph $H$, we can construct an equivalent graph from the union join graph $\mathcal{U}(H)$ which can be built in $\mathcal{O}(|H|+|\mathcal{U}(H)|)$, lowering the overall time complexity of enumeration to $\mathcal{O}(|H|+k)$.
In the rest of this section, we formally introduce the notion of the equivalent graph $G^{\equiv}$. Thereafter, we outline the main algorithm for enumerating join trees, prove its correctness and analyze its complexity. Finally, we consider the case of $\gamma$-acyclic queries.
@@ 4.1 Equivalent Graph
The construction of the equivalent graph $G^{\equiv}$ for a given edge-weighted graph $G$, is based on the following local graph modification, sliding transformation.
@@
Let $G$ be an edge-weighted graph with weight function $\omega$. Given edges $e^{*}=(u,v)$ and $e=(v,w)$ where $\omega(e)<\omega(e^{*})$, sliding $e$ along $e^{*}$ moves $e$’s end node $v$ to $u$ such that $e=(u,w)$.
@@
Figure 2: $T_{W}$ is an MST of the line graph in Figure˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges. $G^{\equiv}$ is the resulting equivalent graph where the thick lines highlight the parallel edges.
We can construct an equivalent graph by repeatedly performing sliding transformations. Sliding may create parallel edges, cycles, and self-loops, so the result is a multigraph.
Definition 8 (Equivalent Graph [12]).
Given weighted graph $G$ with MST $T$ rooted, an equivalent graph $G^{\equiv}$ is obtained from $G$ by exhaustively sliding every edge $e=(v,w)\in E(G)$ along another $e^{*}=(u,v)\in T$ as long as $0pt(T,u)<0pt(T,v)$.
Sliding transformation preserves the identity of each edge $e\in E(G)$ as explained in Section˜4.1. Every spanning tree of $G^{\equiv}$ is an MST of $G$ and vice versa.
Theorem 9 (by Eppstein [12]).
The maximum spanning trees of a weighted graph $G$ are in one-to-one correspondence with the spanning trees of its equivalent graph $G^{\equiv}$. [^2] Minimum spanning trees are considered in the original literature [12].
In the following, we use $p(r)$ to denote the parent of $r$ in a given rooted tree, and $c(r)$ to denote its children. We extend the notation of parent to edges as follows: for any edge $e\in E(T)$, its parent $p(e)\in E(T)$ is another tree edge incident to $e$ and closer to the root $r$. The edges incident to the root do not have a parent and are all siblings. Otherwise, the siblings of $e$ are incident to the same parent, denoted as $s(e)=\left\{e^{\prime}\in E(T)\setminus\{e\}\mid p(e^{\prime})=p(e)\right\}$.
@@
Given the 6-clique line graph $L_{6}$ in Figure˜2(b), we find a rooted MST $T_{W}$ as shown in Figure˜2(a). Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $(T,P)$ whose weight $\omega(T,P)=|\{a\}|=1$ is lighter than its parent tree edge $(W,T)$ with $\omega(W,T)=|\{a,c\}|=2$. We slide along the tree edge $(W,T)$ to the root so that the edge $(T,P)$ becomes $(W,P)$ as shown by the solid curve in Figure˜2(b). Non-tree edges can slide similarly. We consider two examples, $(S,U)$ illustrated with a dashed line and $(Y,T)$ with a dotted line in Figure˜2(a). All other non-tree edges are shown in light gray. We can slide along tree edge $(U,W)$ to the root so that the edge $(S,U)$ becomes $(S,W)$ as shown by the dashed curve in Figure˜2(b). Both ends of $(Y,T)$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in Figure˜2(b). This edge will not appear in any spanning tree of $G^{\equiv}$, and therefore not a part of any MST of $L_{6}$. We refer to such an edge as a non-MST edge, as opposed to an MST edge. By applying sliding transformations to a fixpoint, we obtain an equivalent graph $G^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in Figure˜2(c). For example, the tree edge $(T,P)$ in $T_{W}$ and non-tree edges $(P,U),(P,T),(P,Y)$ are parallel in $G^{\equiv}$ between $P$ and $W$. We can easily verify that each spanning tree of $G^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in Figure˜2(d).
The choice of the initial spanning tree can affect the structure of the equivalent graph, but the order of sliding transformations performed has no impact [12].
@@ 4.2 Enumerating Join Trees of $\alpha$-Acyclic Hypergraphs
The bottleneck of Eppstein’s algorithm for constructing equivalent graphs of general graphs comes from a subroutine that identifies where each edge will eventually slide to. Because each edge can only slide along a heavier edge, it will eventually be “blocked” by a lighter or equally weighted edge along its path to the root. The subroutine essentially performs binary search to find the blocking edge, leading to the $\log n$ factor in the overall complexity. The key to our improvement is to show that for every acyclic hypergraph $H$, we can construct an “equivalent hypergraph” $H^{*}$ that shares the same structure with $H$ and has the same set of join trees, but $H^{*}$ admits a special join tree with monotonically increasing weight from root to leaf, which enables constant-time identification of the blocking edge.
@@
A monotonic weight join tree $\hat{T}\in\mathcal{T}(H)$ is a rooted join tree of $H$ such that for any $e\in\hat{T}$ that has a parent $p(e)$, $\omega(e)>\omega(p(e))$.

### v3

@@ 4 Enumerating Join Trees
Our strategy for enumerating the join trees of a hypergraph $H$ starts from Theorem˜11 which allows us to reduce the problem to the enumeration of maximum spanning trees of the line graph $L(H)$. The best known algorithm for MST enumeration [^3] Eppstein focuses on minimum spanning trees, but the same algorithm applies to maximum spanning trees with flipped comparisons. is due to Eppstein [14] by deriving from the input graph $G$ a so-called equivalent graph $G^{\equiv}$. Every spanning tree of $G^{\equiv}$ corresponds to an MST of $G$ and vice versa. Eppstein gives an algorithm (and proves a matching lower bound) in time $\mathcal{O}(m+n\log n)$ to construct the equivalent graph from an arbitrary weighted graph with $m$ edges and $n$ vertices. Then to enumerate all $k$ MSTs of $G$, he applies existing algorithms to enumerate the spanning trees of $G^{\equiv}$. Since there are optimal spanning tree enumeration algorithms that run in $\mathcal{O}(m+n+k)$ time [24, 37], the overall time complexity to enumerate MSTs is $\mathcal{O}(m+n\log n+k)$. The main result of this section is an algorithm for enumerating join trees leveraging the structure of acyclic hypergraphs and their line graphs. In particular:
  • Given the line graph $L$ of an $\alpha$-acyclic hypergraph, we can construct an equivalent graph of $L$ in $\mathcal{O}(|L|)$, thus enumerating the join trees in $\mathcal{O}(|L|+k)$ time (Theorem˜27).
  • Given any $\gamma$-acyclic hypergraph $H$, we can construct an equivalent graph of $L(H)$ in $\mathcal{O}(|H|)$, lowering the overall time complexity of enumeration to $\mathcal{O}(|H|+k)$ (Theorem˜28).
Note that $|L(H)|$ can be quadratic in $|H|$ while $|H|\in\mathcal{O}(|L(H)|)$, so the second item above yields a better bound.
In the rest of this section, we first define what is an equivalent graph. Then, we present the algorithm for enumerating join trees of $\alpha$-acyclic hypergraphs. Finally, we adapt the algorithm to enumerate join trees of $\gamma$-acyclic hypergraphs more efficiently.
@@ 4.1 Equivalent Graph
The key idea of Eppstein’s algorithm [14] is to construct an equivalent graph $G^{\equiv}$ whose spanning trees one-to-one correspond to the MSTs of the input graph $G$, thereby reducing MST enumeration to spanning tree enumeration.
Definition 15 (Equivalent Graph [14]).
Given a weighted graph $G=((R,E,\rho),\omega)$, a multigraph $G^{\equiv}=((R,E,\rho^{\equiv}),\omega)$ is an equivalent graph of $G$ if the spanning trees of $G^{\equiv}$ one-to-one correspond to the maximum spanning trees of $G$: a set of edges $E^{T}\subseteq E$ induces a spanning tree of $G^{\equiv}$ if and only if $E^{T}$ induces an MST of $G$.
In what follows, we may use the notation $G^{\equiv}(L(H))$ or $G^{\equiv}(H)$ to refer to an equivalent graph of the line graph of hypergraph $H$. Note that $G^{\equiv}$ shares the same vertices and edges of $G$, and they only differ in the incidence function mapping each edge to its endpoints. Such an equivalent graph can be constructed from $G$ by applying a series of sliding transformations: for two incident edges $e,e^{\prime}$ with $\rho(e)=\{r_{1},r_{2}\}$ and $\rho(e^{\prime})=\{r_{2},r_{3}\}$, we can slide $e^{\prime}$ along $e$ by updating $\rho(e^{\prime})=\{r_{1},r_{3}\}$, if $\omega(e^{\prime})<\omega(e)$. Note that the edge $e^{\prime}$ retains its identity after sliding, which explains the need for the incidence function $\rho$. Furthermore, we will use a given rooted MST $T$ as a guide, and only slide edges towards the root. More formally:
@@
Let $T$ be a MST of $G$ rooted at $r$. Let $e$ be an edge such that $\rho(e)=\{r_{1},r_{2}\}$ with $r_{1}$ closer to the root than $r_{2}$ in $T$. If another edge $e^{\prime}$ shares $r_{2}$ with $e$, i.e., $\rho(e^{\prime})=\{r_{2},r_{3}\}$, and $\omega(e^{\prime})<\omega(e)$, then sliding $e^{\prime}$ along $e$ results in a graph $G^{\prime}$ that is identical to $G$, except that the incidence function maps $e^{\prime}$ to $\{r_{1},r_{3}\}$.
The key result by Eppstein shows that applying sliding transformations on $G$ along a rooted MST $T$ to a fixpoint results in an equivalent graph of $G$. To state this formally, we first define an ordering on graphs based on sliding transformations. It is easy to verify the following is a partial order:
Definition 17 (Sliding Partial Order).
We write $G\leq_{T}G^{\prime}$ if $G^{\prime}$ can be obtained from $G$ by applying a sequence of sliding transformations along a rooted MST $T$ of $G$.
Applying sliding transformations to a fixpoint therefore yields a maximal element under the sliding partial order.
Theorem 18 (Sliding produces $G^{\equiv}$ at fixpoint [14]).
Given a weighted graph $G$ and a rooted MST $T$, and let $G^{\equiv}$ be a maximal element under $\leq_{T}$. Then $G^{\equiv}$ is an equivalent graph of $G$.
The choice of the initial rooted MST can affect the structure of the equivalent graph, but the order of sliding transformations performed has no impact [14].
@@
Figure 2: $T_{W}$ is an MST of the line graph in fig.˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges. $G^{\equiv}$ is the resulting equivalent graph where the thick lines highlight the parallel edges.
@@
Given the 6-clique line graph $L_{6}$ in fig.˜2(b), we find a rooted MST $T_{W}$ as shown in fig.˜2(a). Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $\{T,P\}$ whose weight $\omega(\{T,P\})=|\{a\}|=1$ is lighter than its parent tree edge $\{W,T\}$ with $\omega(\{W,T\})=|\{a,c\}|=2$. We slide along the tree edge $\{W,T\}$ to the root so that the edge $\{T,P\}$ becomes $\{W,P\}$ as shown by the solid curve in fig.˜2(b). Non-tree edges can slide similarly. We consider two examples, $\{S,U\}$ illustrated with a dashed line and $\{Y,T\}$ with a dotted line in fig.˜2(a). All other non-tree edges are shown in light gray. We can slide $\{S,U\}$ along tree edge $\{U,W\}$ to become $\{S,W\}$ as shown by the dashed curve in fig.˜2(b). Both ends of $\{Y,T\}$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in fig.˜2(b). This edge will not appear in any spanning tree of $L_{6}^{\equiv}$, and therefore not a part of any MST of $L_{6}$. We refer to such an edge as a non-MST edge, as opposed to an MST edge. Denoting the graph after sliding as $L_{6}^{\prime}$, ordering $L_{6}\leq_{T_{W}}L_{6}^{\prime}$ holds. By applying sliding transformations to a fixpoint, we obtain an equivalent graph $L_{6}^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in fig.˜2(c) (self-loops are omitted). For example, the tree edge $\{T,P\}$ in $T_{W}$ and non-tree edges $\{P,U\},\{P,T\},\{P,Y\}$ become parallel in $L_{6}^{\equiv}$ between $P$ and $W$. Therefore $L_{6}\leq_{T_{W}}L_{6}^{\prime}\leq_{T_{W}}L_{6}^{\equiv}$ holds under the sliding partial order. We can easily verify that each spanning tree of $L_{6}^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in fig.˜2(d).
@@ 4.2 Enumerating Join Trees of $\alpha$-Acyclic Hypergraphs
The bottleneck of Eppstein’s algorithm for constructing equivalent graphs of arbitrary weighted graphs lies in a subroutine that identifies where each edge will eventually slide to. Because each edge can only slide along a heavier edge, it will eventually be “blocked” by a lighter or equally weighted edge along its path to the root. The subroutine essentially performs binary search to find the blocking edge, leading to the $\log n$ factor in the overall complexity. The key to our improvement is to show that for every acyclic hypergraph $H$, we can construct an equivalent hypergraph $H^{*}$ whose join trees one-to-one correspond to those of $H$, but one special join tree of $H^{*}$ has monotonically increasing weight from root to leaf, which enables constant-time identification of the blocking edge.
@@
[…]
We can always construct such a $H^{*}$ given any $\alpha$-acyclic hypergraph $H$:
Theorem 21.
Given an $\alpha$-acyclic hypergraph $H=(X,R,\chi)$, there exists an equivalent hypergraph $H^{*}=(X^{*},R,\chi^{*})$ over the same hyperedge set $R$ that
  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$, and
  • $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
(a) $H^{*}_{6}$
(b) $\hat{T}_{p}$
(c) $\hat{T}_{p}$ with 3 non-tree edges
Figure 3: An equivalent hypergraph $H^{*}_{6}$ and its monotonic weight join tree $\hat{T}_{P}$ rooted at $P$.